OpenStudy (anonymous):

how do u solve y+8=3y-6

OpenStudy (anonymous):

Begin by adding 6 to both sides, so you will have y+14=3y

OpenStudy (anonymous):

then subtract the y from both sides, so you will have 14=2y

OpenStudy (anonymous):

and then you divide both sides by 2 so you get 7=y

OpenStudy (anonymous):

y+8=3y-6 y+3y+8=3y+3-6 4y+8=-6 4y+8-8=-6-8 4y=-14